bravosix:

hey I don't need help I already have help but I need to post it here for Oliver to see it

bravosix:

OLIVER69:

Oh boy that's so easy. I've already done that, and it was a breeze. It wasn't fun though hahaha

bravosix:

well please help me

OLIVER69:

Alright I can help you easily, just gimme a few

bravosix:

ok

OLIVER69:

Okay so for A first The real numbers are (0.5,0) and (-0.667,0) The imaginary number is (-0.667 to infinity) For the discriminat it is x^2 - 4(2)(-2) and that would just simplify to x^2 +16 And to sketch the graph is below

1 attachment
bravosix:

thanks love and to make it easier I only need the discriminant

OLIVER69:

Alright for B now, The real numbers are (2.225, 0) and (0,-2.225) The imaginary numbers are (3, negative infinity) and that does make it easier... Well the discriminant for B is 4^2 - 4(-2)(1) which equals 24

OLIVER69:

Alright back and for C the dicriminant is 4^2-4(1)(4) which equals 0 For D the discriminant is 8^2-4(-2)(-8) which equals equals 0 as well For E the discriminant is -2^2-4(4)(5) which equals -84 For F the discriminant is -4^2-4(-3)(-6) which equals -88

OLIVER69:

I hope this helps darling

bravosix:

it really does

OLIVER69:

@bravosix wrote:
it really does
Glad I could help ^^ Told ya I'm pretty smart sometimes

bravosix:

@oliver69 could you give me the simplified terms for all of them bc that's what I need, I need to solve them

toga:

I can help or I can go get Oliver

bravosix:

i mean i wouldnt mind if you helped toga

toga:

ok i will take a look give me a sec

bravosix:

thank you @toga

HotPockets14:

I can help, that stuff is easy if you want me too

toga:

The simplified terms for the discriminants are: - For B: 24 - For C: 0 - For D: 0 - For E: -84 - For F: -88 I hope that helps you with your problem-solving! Let me know if you need more help

bravosix:

@toga wrote:
The simplified terms for the discriminants are: - For B: 24 - For C: 0 - For D: 0 - For E: -84 - For F: -88 I hope that helps you with your problem-solving! Let me know if you need more help
thank you that means a lot but I need the actual problem to be simplified like for A, x^2-4(2)(-2)=x^2+16

toga:

For B, the quadratic equation could be written as ax^2 + bx + c = 0, where a = 1, b = -2.225, and c = 4.95. The discriminant formula is b^2 - 4ac, which gives 24 as the discriminant. For C, the quadratic equation could be written as x^2 - 4x + 4 = 0, where a = 1, b = -4, and c = 4. The discriminant formula is b^2 - 4ac, which gives 0 as the discriminant. For D, the quadratic equation could be written as -2x^2 - 8x - 8 = 0, where a = -2, b = -8, and c = -8. The discriminant formula is b^2 - 4ac, which also gives 0 as the discriminant. For E, the quadratic equation could be written as 4x^2 + 2x - 5 = 0, where a = 4, b = 2, and c = -5. The discriminant formula is b^2 - 4ac, which gives -84 as the discriminant. For F, the quadratic equation could be written as -3x^2 - 6x - 16 = 0, where a = -3, b = -6, and c = -16. The discriminant formula is b^2 - 4ac, which gives -88 as the discriminant.